OpenStudy (anonymous):

integrate of 2x^3 / x^2-4?

OpenStudy (mimi_x3):

\[\int\limits\frac{2x^{3}}{x^{2}-4} dx\]?

OpenStudy (anonymous):

yup. :)

OpenStudy (anonymous):

I have a correct answer but i dont have the solution. can you show me?

OpenStudy (mimi_x3):

\[2\int\limits\frac{x^{3}}{x^{2}-4} dx\] looks like a long division..

OpenStudy (shubhamsrg):

let x^2 -4 = t thus dt = 2xdx now 2x^3 dx= (2xdx)(x^2) = (t+4)dt so we are left with (t+4)/t dt try now..

OpenStudy (anonymous):

\[\int\limits {} \frac{x(x^{2}-4)+4x}{x^{2}-4}\]

OpenStudy (anonymous):

\[2\int\limits x+\frac{4x}{x^{2}-4}\]

OpenStudy (anonymous):

do you understand now???

OpenStudy (anonymous):

\[x2+4\ln |x^{2} - 4|+C\]

OpenStudy (anonymous):

that's the true answer.

OpenStudy (anonymous):

\[2[\frac{x^{2}}{2}]\] is the integral og the first term..... for second term take \[x^{2}-4=t\] \[2xdx=dt\]

OpenStudy (anonymous):

\[\int\limits \frac{4xdx}{x^{2}-4}=\int\limits \frac{2dt}{t}\]

OpenStudy (anonymous):

\[=2\ln(t)\]

OpenStudy (anonymous):

\[t=x^{2}-4\] adding first and the second \[2[\frac{x^{2}}{2}+2\ln|x^{2}-4|]+c\] \[x^{2}+4\ln|x^{2}-4|+c\]